# ConceptPose: Training-Free Zero-Shot Object Pose Estimation using Concept Vectors

### v2

@@ 4.5 Ablation Studies and Analysis
Extended Experiemnt on Few-shot Pose Tracking. Following FoundationPose [foundationpose2024] and UA-Pose [uapose2025], we present an additional evaluation of few-shot pose tracking on the test split of YCB-Video filtered to BOP test targets. While our pipeline was not specifically designed for this use case, we evaluate it to demonstrate the flexibility of our concept-based approach and its generalizability to related tasks. For each object instance within a scene, we select 2 reference frames using icosphere-based farthest point sampling on the viewing sphere to maximize viewpoint diversity, matching FoundationPose’s reference selection strategy. These reference frames are used to build a instance-level concept model via multi-view aggregation of semantic saliency maps. We then evaluate pose estimation on all remaining frames within the same scene, excluding the reference frames from the test set. In total we evaluate 4068 A-Q pairs. Table 4 shows few-shot tracking results on YCB-Video. ConceptPose achieves 90.1% ADD-AUC and 95.4% ADD-S-AUC using 2 static reference frames, outperforming FoundationPose (87.4%/94.3%) without training. While UA-Pose achieves higher accuracy (92.8%/96.5%) due to online object completion that progressively refines the model during testing, both ConceptPose and FoundationPose use fixed models throughout. Our 2.7-point ADD-AUC gain over FoundationPose demonstrates that concept vectors enable competitive few-shot performance across large viewpoint variations.
Limitations. ConceptPose inherits computational characteristics shared with other training-free VLM-based approaches [foundationpose2024, sam6d2024]. Runtime remains approximately 7 seconds per image pair (6.8 seconds with voxelization), stemming primarily from VLM inference overhead—a limitation shared across methods leveraging dense vision-language models. As VLM architectures advance, we expect proportional runtime improvements. Performance degrades under extreme viewpoint changes for highly asymmetric objects and severe occlusions, challenges common to correspondence-based methods [foundationpose2024, gen6d2022]. A promising direction is extending ConceptPose to category-level training-free pose estimation requiring no reference view at all—leveraging concept vectors’ inherent category-level nature to estimate poses directly from category names, eliminating even the one-shot requirement.
@@
FoundationPose [foundationpose2024] 87.4 94.3
UA-Pose [uapose2025] 92.8 96.5
@@ 5 Conclusion
We presented ConceptPose, a training-free and model-free approach for 6D object pose estimation that leverages VLM explainability for language-driven semantic reasoning. By querying an LLM to generate concept labels and using VLM saliency maps for spatial localization, we construct dense 3D concept vectors that enable correspondence-based pose estimation without learned matching networks. Our method achieves state-of-the-art results on REAL275 [wang2019] (71.5% ADD(-S), +33.6%), TYOL [bop_2018] (55.0%, +59.0%), and YCB-Video [xiang2018] (41.2%, +82.3%), significantly outperforming existing methods including those requiring training—demonstrating that conceptual reasoning can surpass learned geometric features. ConceptPose also demonstrates strong few-shot tracking performance (90.1% ADD-AUC), surpassing FoundationPose [foundationpose2024] without training.
@@ 6.1 Backbone Ablation
To demonstrate ConceptPose’s adaptability across different vision-language architectures, we evaluate five backbones on REAL275 (Table 5): three SigLIP2 [siglip2023] variants (giant-384, large-384, base-384), CLIP ViT-L/14@336px [clip_radford2021], and DINOv3-L/16 [simeoni2025dinov3] with the dinotxt text grounding head [jose2024dinotxt]. Our baseline method is SigLIP2-giant-384. All experiments maintain identical pipeline configurations without voxelization to isolate the impact of backbone architecture.
SigLIP2 and CLIP (GradCAM-based). For SigLIP2-giant-384 (our baseline), SigLIP2-large-384, and SigLIP2-base-384, we use Hugging Face implementations with GradCAM [selvaraju2020] applied to the post_layernorm layer. All SigLIP2 variants use $\times384$ input resolution. For CLIP ViT-L/14@336px, we target visual.transformer.resblocks[-1].ln_1 (layer norm of the final transformer block) at its native $\times336$ resolution. Both methods compute image-text similarity through normalized dot products, with GradCAM extracting gradient-weighted spatial activations from intermediate transformer features.
@@ 8 Extended Qualitative Results
Figures 6 and 7 present 10 extended qualitative visualizations of ConceptPose across all four evaluation datasets (REAL275, Toyota-Light, YCB-Video, LINEMOD), with saliency maps and estimated poses. It demonstrates ConceptPose’s ability to generate semantically meaningful concept activations and accurate pose estimates across diverse object categories and imaging conditions.

### v3

@@ 4.5 Ablation Studies and Analysis
Extended Experiment on Few-shot Pose Tracking. Following FoundationPose [53] and UA-Pose [26], we present an additional evaluation of few-shot pose tracking on the test split of YCB-Video filtered to BOP test targets. While our pipeline was not specifically designed for this use case, we evaluate it to demonstrate the flexibility of our concept-based approach and its generalizability to related tasks. For each object instance within a scene, we select 2 reference frames using icosphere-based farthest point sampling on the viewing sphere to maximize viewpoint diversity, matching FoundationPose’s reference selection strategy. These reference frames are used to build a instance-level concept model via multi-view aggregation of semantic saliency maps. We then evaluate pose estimation on all remaining frames within the same scene, excluding the reference frames from the test set. In total we evaluate 4068 A-Q pairs. Table 4 shows few-shot tracking results on YCB-Video. ConceptPose achieves 90.1% ADD-AUC and 95.4% ADD-S-AUC using 2 static reference frames, outperforming FoundationPose (87.4%/94.3%) without training. While UA-Pose achieves higher accuracy (92.8%/96.5%) due to online object completion that progressively refines the model during testing, both ConceptPose and FoundationPose use fixed models throughout. Our 2.7-point ADD-AUC gain over FoundationPose demonstrates that concept vectors enable competitive few-shot performance across large viewpoint variations.
Limitations. ConceptPose inherits computational characteristics shared with other training-free VLM-based approaches [53, 29]. Runtime remains approximately 7 seconds per image pair (6.8 seconds with voxelization), stemming primarily from VLM inference overhead—a limitation shared across methods leveraging dense vision-language models. As VLM architectures advance, we expect proportional runtime improvements. Performance degrades under extreme viewpoint changes for highly asymmetric objects and severe occlusions, challenges common to correspondence-based methods [53, 33]. A promising direction is extending ConceptPose to category-level training-free pose estimation requiring no reference view at all—leveraging concept vectors’ inherent category-level nature to estimate poses directly from category names, eliminating even the one-shot requirement.
@@
FoundationPose [53] 87.4 94.3
UA-Pose [26] 92.8 96.5
@@ 5 Conclusion
We presented ConceptPose, a training-free and model-free approach for 6D object pose estimation that leverages VLM explainability for language-driven semantic reasoning. By querying an LLM to generate concept labels and using VLM saliency maps for spatial localization, we construct dense 3D concept vectors that enable correspondence-based pose estimation without learned matching networks. Our method achieves state-of-the-art results on REAL275 [51] (71.5% ADD(-S), +33.6%), TYOL [18] (55.0%, +59.0%), and YCB-Video [55] (41.2%, +82.3%), significantly outperforming existing methods including those requiring training—demonstrating that conceptual reasoning can surpass learned geometric features. ConceptPose also demonstrates strong few-shot tracking performance (90.1% ADD-AUC), surpassing FoundationPose [53] without training.
@@ 6.1 Backbone Ablation
To demonstrate ConceptPose’s adaptability across different vision-language architectures, we evaluate five backbones on REAL275 (Table 5): three SigLIP2 [48] variants (giant-384, large-384, base-384), CLIP ViT-L/14@336px [39], and DINOv3-L/16 [42] with the dinotxt text grounding head [21]. Our baseline method is SigLIP2-giant-384. All experiments maintain identical pipeline configurations without voxelization to isolate the impact of backbone architecture.
SigLIP2 and CLIP (GradCAM-based). For SigLIP2-giant-384 (our baseline), SigLIP2-large-384, and SigLIP2-base-384, we use Hugging Face implementations with GradCAM [41] applied to the post_layernorm layer. All SigLIP2 variants use $\times384$ input resolution. For CLIP ViT-L/14@336px, we target visual.transformer.resblocks[-1].ln_1 (layer norm of the final transformer block) at its native $\times336$ resolution. Both methods compute image-text similarity through normalized dot products, with GradCAM extracting gradient-weighted spatial activations from intermediate transformer features.
@@ 7 Performance Analysis by Object
8 Continuous Concept Vectors vs. Binary Masks
Our GradCAM-based saliency provides continuous activations, as opposed to binary masks produced by segmentation-based approaches. To validate the advantage of soft activations, we create binary concept vector masks by thresholding at 0.5 and evaluate on REAL275. As shown in Table 10, both ADD(-S) (71.5% $\rightarrow$ 30.5%) and BOP AR (60.4% $\rightarrow$ 28.2%) decrease strongly, indicating that continuous concept vectors are important for establishing accurate correspondences.
Table 10: Continuous concept vectors vs. binary masks (thresholded at 0.5) on REAL275.
ADD(-S) BOP AR ADD ADD-S $10^{\circ}$/5cm $5^{\circ}$/2cm
Concept vectors 71.5 60.4 60.6 90.8 47.2 26.0
Binary masks (thr.=0.5) 30.5 28.2 25.2 62.7 17.7 6.1
9 Effect of Concept Quantity on Correspondence Quality
Figure 6: Pose estimation with increasing number of concepts (3, 10, 15) on a car object.
Spatially extended and repeated parts are a strength of our method: a single spatially localized concept (e.g., car door) can anchor the object, but cannot resolve its scale or orientation. In contrast, a repeated concept (e.g., wheel) helps resolve the object’s location, scale, and one additional DoF of orientation. This is analogous to a 2D line segment resolving 4 DoF while a point resolves only 2 DoF. Figure 6 shows this behavior with increasing number of concepts (3, 10, 15) on a car object.
10 Occlusion Robustness Analysis
We present a correlation analysis between the amount of occlusion and pose error (ADD-S) on all objects in LINEMOD-Occlusion (LM-O), sampling 20 anchor-query pairs randomly for each object. Figure 7 shows the results. Regression analysis shows $R^{2}<0.3$, indicating no linear correlation between occlusion and error. ConceptPose handles occlusion through several design choices: we crop the input image to the object’s bounding box before VLM inference to alleviate potential contamination due to the patch nature of ViTs, backproject only pixels within the object mask to 3D to reduce background noise, and utilize RANSAC’s robustness against partial observations by finding the largest geometrically consistent subset among visible points.
Figure 7: Correlation between occlusion ratio and ADD-S error on LINEMOD-Occlusion. Each subplot shows one object with 20 randomly sampled anchor-query pairs. Regression analysis shows $R^{2}<0.3$, indicating no linear correlation between occlusion and error.
11 Cross-Instance Concept Generalization
In our task, concepts are inherently shared across instances within a category—not per instance. Figure 8 visualizes the same concepts (lens, grip, controls, hot_shoe) activated across different camera instances, demonstrating our training-free approach’s ability to generalize across instances, unlike learned geometric descriptors.
Figure 8: Cross-instance concept activation on different camera instances. The same category-level concepts [lens, grip, controls, hot_shoe] are consistently localized across diverse instances.
@@ 12 Extended Qualitative Results
Figures 9 and 10 present 10 extended qualitative visualizations of ConceptPose across all four evaluation datasets (REAL275, Toyota-Light, YCB-Video, LINEMOD), with saliency maps and estimated poses. It demonstrates ConceptPose’s ability to generate semantically meaningful concept activations and accurate pose estimates across diverse object categories and imaging conditions.